# Tülu 3: Pushing Frontiers in Open Language Model Post-Training

### v2

@@ Abstract
00footnotetext: Core contributors.Language model post-training is applied to refine behaviors and unlock new skills across a wide range of recent language models, but open recipes for applying these techniques lag behind proprietary ones. The underlying training data and recipes for post-training are simultaneously the most important pieces of the puzzle and the portion with the least transparency. To bridge this gap, we introduce Tülu 3, a family of fully-open state-of-the-art post-trained models, alongside its data, code, and training recipes, serving as a comprehensive guide for modern post-training techniques. Tülu 3, which builds on Llama 3.1 base models, achieves results surpassing the instruct versions of Llama 3.1, Qwen 2.5, Mistral, and even closed models such as GPT-4o-mini and Claude 3.5-Haiku. The training algorithms for our models include supervised finetuning (SFT), Direct Preference Optimization (DPO), and a novel method we call Reinforcement Learning with Verifiable Rewards (RLVR). With Tülu 3, we build a multi-task evaluation scheme for post-training with development and unseen evaluations, standard benchmark implementations, and substantial decontamination of existing open datasets on said benchmarks. We conclude with analysis and discussion of training methods that did not reliably improve performance.
@@
Tülu 3 70B | https://hf.co/allenai/Llama-3.1-Tulu-3-70B |
Tülu 3 8B | https://hf.co/allenai/Llama-3.1-Tulu-3-8B |
Tülu 3 Data | https://hf.co/collections/allenai/tulu-3-datasets- |
673b8df14442393f7213f372 | ||
Tülu 3 Code | https://github.com/allenai/open-instruct |
Tülu 3 Eval | https://github.com/allenai/olmes |
Demo | https://playground.allenai.org/ |
@@
Python Code | allenai/tulu-3-sft-personas-code |
@@ 1 Introduction
To close the gap between open and closed post training, we introduce Tülu 3, a family of open state-of-the-art post-trained models, alongside all of the data, training recipes, code, infrastructure, and evaluation framework. Integrating partial details from proprietary methods with novel techniques and established academic research, Tülu 3 pushes the boundaries of research in post-training. The advancements of Tülu 3 are attributed to Tülu 3 Data, new permissively licensed training datasets targeting core skills, Tülu 3 Eval, evaluation suite and tools to establish clear performance goals and guide improvement through training stages, and Tülu 3 Recipe, an advanced multi-stage training pipeline incorporating new algorithmic advancements in reinforcement learning, cutting-edge infrastructure, and rigorous experimentation to optimize data mixes, methods, and parameters across various training stages.
@@
Our best performing recipe yields Tülu 3 models that outperform the state-of-the-art post-trained open-weight models of the same size such as Llama 3.1 Instruct (dubey2024llama), Qwen2.5 Instruct (qwen2.5), or Mistral-Instruct (mistral2024ministraux), and at the large 70B size Tülu matches the offerings of closed providers such as Claude 3.5 Haiku and GPT-4o mini.
@@
Reinforcement learning with verifiable rewards, an RL-based method that only gets a reward if the model’s completions are verified to be correct,
Advanced infrastructure, details, and code to facilitate the successful implementation of large models

### v3

@@ Abstract
Language model post-training is applied to refine behaviors and unlock new skills across a wide range of language models, but open recipes for applying these techniques lag behind proprietary ones. The underlying training data and recipes for post-training are simultaneously the most important pieces of the puzzle and the portion with the least transparency. To bridge this gap, we introduce Tülu 3, a family of fully-open state-of-the-art post-trained models, alongside its data, code, and training recipes, serving as a comprehensive guide for modern post-training techniques. Tülu 3, which builds on Llama 3.1 base models, achieves results surpassing the instruct versions of Llama 3.1, Qwen 2.5, Mistral, and even closed models such as GPT-4o-mini and Claude 3.5-Haiku. The training algorithms for our models include supervised finetuning (SFT), Direct Preference Optimization (DPO), and a novel method we call Reinforcement Learning with Verifiable Rewards (RLVR). With Tülu 3, we build a multi-task evaluation scheme for post-training with development and unseen evaluations, standard benchmark implementations, and substantial decontamination of existing open datasets on said benchmarks. We conclude with analysis and discussion of training methods that did not reliably improve performance.
@@
Tulu 3 8B:]Llama-3.1-Tulu-3-8B
[…]
Tulu 3 70B:]Llama-3.1-Tulu-3-70B
[…]
Tulu 3 405B:]Llama-3.1-Tulu-3-405B
[…]
Tulu 3 Data:]tulu-3-datasets-673b8df14442393f7213f372
[…]
Tulu 3 Code:]open-instruct
[…]
Tülu 3 Eval:]olmes
[…]
Demo:]playground.allenai.org
@@
| Stage | Llama 3.1 405B | |
| Base Model | meta-llama/Llama-3.1-405B | |
| SFT | allenai/Llama-3.1-Tulu-3-405B-SFT | |
| Final Model | ||
| (DPO) | allenai/Llama-3.1-Tulu-3-405B |
@@
| Python Code | allenai/tulu-3-sft-personas-code | |
@@ 1 Introduction
To close the gap between open and closed post training, we introduce Tülu 111A tülu is a hybrid camel bred between Bactrian camel and dromedary: https://en.wikipedia.org/wiki/Hybrid_camel. 3, a family of open state-of-the-art post-trained models, alongside all of the data, training recipes, code, infrastructure, and evaluation framework. Integrating partial details from proprietary methods with novel techniques and established academic research, Tülu 3 pushes the boundaries of research in post-training. The advancements of Tülu 3 are attributed to Tülu 3 Data, new permissively licensed training datasets targeting core skills, Tülu 3 Eval, an evaluation suite and tools to establish clear performance goals and guide improvement through training stages, and Tülu 3 Recipe, an advanced multi-stage training pipeline incorporating new algorithmic advancements in reinforcement learning, cutting-edge infrastructure, and rigorous experimentation to optimize data mixes, methods, and parameters across various training stages.
@@
Our best performing recipe yields Tülu 3 models that outperform the state-of-the-art post-trained open-weight models of the same size such as Llama 3.1 Instruct (dubey2024llama) or Mistral-Instruct (mistral2024ministraux), and at the large 70B size Tülu matches the offerings of closed providers such as Claude 3.5 Haiku and GPT-4o mini. Furthermore, at 405B size our model performs competitively against DeepSeek v3 (deepseekai2024deepseekv3technicalreport) and GPT 4o (11-24).
@@
Reinforcement learning with verifiable rewards, an RL-based method that only gets a reward if the model’s completions are verified to be correct, and
Advanced infrastructure, details, and code to facilitate the successful implementation of large models.